if the area of a triangle is 60 sq m and the base is 20 m, what is the altitude h. ISBN 0-13-073784-4
Answer by nerdybill(7384)

if the area of a triangle is 60 sq m and the base is 20 m, what is the altitude h.
.
For any triangle,
area = (1/2)bh
where
b is length of base
h is height
.
Plug in what was given into:
area = (1/2)bh
60 = (1/2)(20)h
solve for h:
60 = (1/2)(20)h
60 = (10)h
6 m = h (altitude)